OpenStudy (anonymous):

How would i find this answer, 8= √2^(4x+8)?

OpenStudy (raden):

convert 8=2^3 and √(2^(4x+8)) = 2^((4x+8)/2) = 2^(2x+4)

OpenStudy (raden):

just take the power of both, then solve for x

OpenStudy (anonymous):

i have no idea what you just did

OpenStudy (raden):

well, look this 8= √(2^(4x+8)) 2^3 = 2^((4x+8)/2) 2^3 = 2^(2x+4) 3 = 2x+4 solve for x from here

OpenStudy (anonymous):

i got -2 but it equals zero

OpenStudy (raden):

still incorrect

OpenStudy (raden):

3 = 2x+4 or 2x+4 = 3 susbtracted by 4 to both sides, what u get

OpenStudy (anonymous):

i get -1=2x

OpenStudy (raden):

nice... then divided by 2 to both sides, what u get

OpenStudy (anonymous):

i divided 2 by -1 and got -2

OpenStudy (raden):

nopes... -1 = 2x or 2x = -1 divided by 2, both sides it means 2x/2 = -1/2 x = - 1/2

OpenStudy (raden):

so, the answer is x=-1/2

OpenStudy (anonymous):

oh, can you explain the first process above in words?

OpenStudy (raden):

which parts ?

OpenStudy (anonymous):

8= √(2^(4x+8)) 2^3 = 2^((4x+8)/2) 2^3 = 2^(2x+4) 3 = 2x+4 this part

OpenStudy (raden):

ok.. first 8 = 2^3, because 2^3 = 2*2*2 = 8 right ?

OpenStudy (raden):

hello ?

OpenStudy (anonymous):

yes

OpenStudy (raden):

ok, next... for √(2^(4x+8)), u have to go out from squareroot, use the property :